# Canvil: Designerly Adaptation for LLM-Powered User Experiences

### v1

@@ 9. Conclusion
In this paper, we introduced the practice of designerly adaptation to encourage designers to engage with LLMs as an adaptable design material when crafting LLM-powered user experiences. We first built up a characterization of designerly adaptation through a formative interview study with 12 designers. We then developed Canvil, a Figma widget that operationalizes designerly adaptation by enabling designers to iteratively author, test, and share system prompts within Figma’s collaborative design environment. We used Canvil as a technology probe to explore the integration of designerly adaptation in UX practice through a group-based design study with 17 designers. We observed numerous promising approaches taken by designers such as iteratively tinkering with different adaptation strategies, sensitizing the model to nuances in user needs, and reasoning about interface affordances using research-informed model behavior; these approaches’ promises were amplified once adaptation was embraced as a collaborative practice. Our work illuminates paths for new processes and tools to spotlight designers’ user-centered perspectives and expertise for more responsible and thoughtful deployment of LLM-powered technologies.
@@ Appendix A Formative Study Participants
P1
Product Designer
6–10
Domain-specific Data QA
Visual/Industrial Design; Computing
Denmark
P2
Interaction Designer
6–10
Conversational Search
Visual/Industrial Design
Canada
P3
Principal Product Designer
6–10
Domain-specific Data QA
Visual/Industrial Design
U.S.
P4
Content Designer
6–10
Conversational Search; Domain-specific Data QA
Humanities
U.S.
P5
Principal Content Designer
6–10
Conversational Search
Humanities
U.S.
P6
Content Designer
1–2
Recommendation
Humanities
U.S.
P7
UX Researcher
3–5
Domain-specific Data QA
Visual/Industrial Design; Social & Behavioral Sciences
Ireland
P8
Product Designer
3–5
Creativity Support Tools
Visual/Industrial Design
U.S.
P9
Senior Designer
11+
Conversational Search; Creativity Support Tools
Visual/Industrial Design
U.S.
P10
UX Researcher
11+
Domain-specific Data QA
Social & Behavioral Sciences
U.S.
P11
Product Designer
3–5
Text Editing & Generation; Domain-specific Data QA
Social & Behavioral Sciences; Computing
U.S.
P12
Principal Content Design Manager
11+
Text Editing & Generation; Domain-specific Data QA
Humanities
U.S.
Table 1. Details of our participants (job title, years of experience in design, LLM application area of their product/feature, and education background before starting their current role) in our formative study.
@@ Appendix B Canvil System Prompt Template
[Example inputs/outputs]
@@
<|im_end|>
Appendix C Design Study Participants
@@
G1
P1
No
Design Student
3–5
Visual/Industrial Design
P2
No
Design Student
1–2
Visual/Industrial Design
P3
No
Product Designer
3–5
Visual/Industrial Design; Computing; Information
G2
P4
No
Product Designer
1–2
Visual/Industrial Design; Computing
P5
No
Product Designer
3–5
Visual/Industrial Design
P6
No
UX Designer
1–2
Visual/Industrial Design
G3
P7
No
Digital Designer
1–2
Visual/Industrial Design
P8
No
Accessibility & Product Designer
1–2
Visual/Industrial Design; Management
P9
No
Design Student
1–2
Visual/Industrial Design; Computing; Information
G4
P10
Yes
Principal Content Design Manager
11+
Humanities
P11
Yes
Principal Content Designer
6–10
Humanities, Visual/Industrial Design; Social & Behavioral Sciences
G5
P12
Yes
Senior Product Designer
3–5
Visual/Industrial Design
P13
Yes
Product Designer
3–5
Visual/Industrial Design
P14
Yes
Product Designer
3–5
Social & Behavioral Sciences; Computing
G6
P15
Yes
Content Designer
1–2
Humanities
P16
Yes
UX Researcher
11+
Social & Behavioral Sciences
P17
Yes
UX/UI Designer
1–2
Information
Table 2. Details of our participants (job title, years of experience in design, LLM application area of their product/feature, and educational background before starting their current role) in our design study. All participants were based in the U.S..
Appendix D Feasto User Groups
D.1. Dimensions of Variance
@@
D.2. User Group Descriptions
D.2.1. North America

### v2

@@ 9. Conclusion
As LLMs become increasingly embedded in our everyday applications, designers are empowered to craft effective LLM-powered UX. Through interviews with 12 designers, we identified a need for a design code that can facilitate a two-way translation between LLM behavior and design requirements. We proposed designerly adaptation as such a design code. We then developed Canvil, a Figma widget that operationalizes designerly adaptation by enabling designers to iteratively author, tinker with, and share adapted LLMs as a novel design material within Figma’s collaborative canvas environment. We used Canvil as a technology probe to explore the integration of designerly adaptation in UX practice through a group-based design study with 17 designers in 6 groups. Designers effectively made use of the two-way translation between design requirements and their UX designs by using LLMs adapted with their requirements to improve interface affordances, while also using their designs to define additional requirements for model behavior; these approaches’ promises were amplified once designerly adaptation was embraced as a collaborative practice. Our work illuminates paths for new processes and tools to spotlight designers’ user-centered expertise for more responsible and thoughtful deployment of LLM-powered technologies.
@@ Appendix A Formative Study Participants
P1 UX Designer 6–10 Domain-specific QA Visual/Industrial Design; Computing Denmark
P2 UX Designer 6–10 Conversational Search Visual/Industrial Design Canada
P3 Principal UX Designer 6–10 Domain-specific QA Visual/Industrial Design U.S.
P4 Content Designer 6–10 Conversational Search; Domain-specific QA Humanities U.S.
P5 Principal Content Designer 6–10 Conversational Search Humanities U.S.
P6 Content Designer 1–2 Recommendation Humanities U.S.
P7 UX Researcher 3–5 Domain-specific QA Visual/Industrial Design; Social & Behavioral Sciences Ireland
P8 UX Designer 3–5 Creativity Support Tools Visual/Industrial Design U.S.
P9 Senior UX Designer 11+ Conversational Search; Creativity Support Tools Visual/Industrial Design U.S.
P10 UX Researcher 11+ Domain-specific QA Social & Behavioral Sciences U.S.
P11 UX Designer 3–5 Text Editing & Generation; Domain-specific QA Social & Behavioral Sciences; Computing U.S.
P12 Principal Content Designer 11+ Text Editing & Generation; Domain-specific QA Humanities U.S.
Table 1. Details of our participants (job title, years of experience in design, LLM application area of their product/feature, and education background before starting their current role) in our formative study. All participants used Figma in their day-to-day work, and have experience designing user experiences for LLM-powered products.
[…]
A table summarizing the background information of 12 participants (P1 to P12) involved in our formative study, categorized by job title, years of experience (YoE), LLM application area, educational background, and region. Column 1: Participant Number (P#): Lists participant identifiers from P1 to P12. Column 2: Job Title: Includes titles such as UX Designer, Content Designer, Principal UX Designer, Principal Content Designer, and UX Researcher. Column 3: Years of Experience (YoE): Ranges from 1–2 years to 11+ years. Column 4: LLM Application Area: Lists areas such as Domain-specific QA, Conversational Search, Text Editing & Generation, Recommendation, Creativity Support Tools, and Domain-specific QA. Column 5: Educational Background: Includes fields like Visual/Industrial Design, Humanities, Computing, Social Behavioral Sciences. Column 6: Region: Specifies the region of the participants, including U.S., Canada, Denmark, and Ireland.
@@ Appendix B Canvil System Prompt Template
{Example inputs/outputs pairs}
@@
<|im\_end|>
Appendix C Design Study Recruitment Details
We recruited our participants from two channels. First, we distributed study invites to designers via email, professional interest groups, and word of mouth at a technology company. Some of these invitees took part in our formative study. We also snowball sampled by asking our invitees to refer us to other designers they work with who may also be interested in participating. From this channel, we only selected designers with prior experience working on LLM-powered products and features. Second, we sent study invites to a Slack workspace for HCI and design maintained by a large, public university in the United States. The population in the Slack consists primarily of design students and early-career designers. From this channel, we only selected those without any experience working on LLM-powered products and features. Our goal of recruiting from these two channels was to capture potential disparities that may arise in our findings hinging on prior experience working with LLMs. In the end, however, we did not notice meaningful qualitative differences in results between the two groups, except that those with LLM design experience drew more connections between adaptation and their past workflows for tinkering with LLMs or AI in general.
We organized all candidates from the technology company into groups based on member preferences and availabilities, and did the same for candidates from the academic institution. In the end, we had 8 participants from the technology company and 9 from the academic institution.
Appendix D Design Study Participants
See Table 2.
@@ Appendix D Design Study Participants
G1 P1 No Design Student 3–5 Visual/Industrial Design
P2 No Design Student 1–2 Visual/Industrial Design
P3 No UX Designer 3–5 Visual/Industrial Design; Computing; Information
G2 P4 No UX Designer 1–2 Visual/Industrial Design; Computing
P5 No UX Designer 3–5 Visual/Industrial Design
P6 No UX Designer 1–2 Visual/Industrial Design
G3 P7 No UX Designer 1–2 Visual/Industrial Design
P8 No UX Designer 1–2 Visual/Industrial Design; Management
P9 No Design Student 1–2 Visual/Industrial Design; Computing; Information
G4 P10 Yes Principal Content Designer 11+ Humanities
P11 Yes Principal Content Designer 6–10 Humanities, Visual/Industrial Design; Social & Behavioral Sciences
G5 P12 Yes Senior UX Designer 3–5 Visual/Industrial Design
P13 Yes UX Designer 3–5 Visual/Industrial Design
P14 Yes UX Designer 3–5 Social & Behavioral Sciences; Computing
G6 P15 Yes Content Designer 1–2 Humanities
P16 Yes UX Researcher 11+ Social & Behavioral Sciences
P17 Yes UX Designer 1–2 Information
Table 2. Details of our participants (job title, years of experience in design, LLM application area of their product/feature, and educational background before starting their current role) in our design study. All participants were based in the U.S., and used Figma in their day-to-day work.
[…]
A table summarizing the background information of participants (P1 to P17) grouped into six groups (G1 to G6) in our design study, based on their prior LLM design experience, job title, years of experience (YoE), and educational background. Column 1: Group Number (G#): Participants are grouped from G1 to G6. Column 2: Participant Number (P#): Lists participant identifiers from P1 to P17. Column 3: Prior LLM Design Experience?: Indicates whether participants had prior experience designing LLMs (”Yes” or ”No”). Column 4: Job Title: Includes titles like Design Student, UX Designer, Senior UX Designer, Content Designer, and Principal Content Designer. Column 5: Years of Experience (YoE): Ranges from 1–2 years to 11+ years. Column 6: Educational Background: Lists fields such as Visual/Industrial Design, Computing, Information, Social & Behavioral Sciences, Humanities, and Management.
Appendix E Feasto User Groups
E.1. Dimensions of Variance
@@
E.2. User Group Descriptions
E.2.1. North America